Exercise · ZW-BERNOULLI-DISTRIBUTION-004 · ★☆☆ Basic

A fair die is rolled and \(X=1\) if a 6 shows, \(X=0\) otherwise. Compute \(\operatorname{Var}(X)\).

Choose one option.

Show answer
  • B. \(\tfrac{5}{36}\)

Explanation

Here \(p=\tfrac16\), so \(p(1-p)=\tfrac16\cdot\tfrac56=\tfrac5{36}\).

Study Probability and statistics in ZWMath →

Review the concept: Bernoulli distribution

A Bernoulli variable \(X\) only takes the values 1 (success) and 0 (failure); its only parameter \(p\) is the probability \(P\) of success.

\[P(X=1)=p,\quad P(X=0)=1-p\]

More practice problems: bernoulli distribution

← ZW-BERNOULLI-DISTRIBUTION-003 ZW-BERNOULLI-DISTRIBUTION-005 →